Grade 12Physical Chemistry

Photons of an electromagnetic radiation of wavelength 62 nm strikes a metal sheet having work function 10.625 eV. The minimum possible wavelength of emitted electrons is

Sujit Kumar
7 Years agoGrade 12

1 Answer

Brij
7 Years ago
No emission will occur because
As we calculate threshold wavelength of emitted electron 
E°=hc/^(lamda)°
^(lamda)°=hc/10.625×1.6×10^-19
(^)Lamda°=6.6×10^-34×1.6×10^-19/10.625
^(lamda)°=186nm 
This is the minimum wavelength of emitted electron..